How do I add these numbers together and divide by 2 to find a real number?
7/8 and 9/10

Found 2 solutions by stanbon, rcmcc:
Answer by stanbon(75887)

How do I add these numbers together and divide by 2 to find a real number?
7/8 and 9/10
-----------
Get the same denominator on both fraction:
The least common multiple of 8 and 10 is 40
Rewrite each fraction with 40 as its denominator:
7*5/8*5 + 9*4/10*4
=35/40 + 36/40
=71/40
Now, divide by 2:
[71/40]/2
=71/[40*2]
=71/80
Cheers,
stan H.

Answer by rcmcc(152)

9/10 and 7/8
to add these items together you must find a comon denominator in this case an easy denominator is 80,
multiply 9/10 by 8 to get 72/80
multiply 7/8 by 10 to get 70/80
now you can add them together to get 142/80 you can divide the top by 2 to find the answer
142/2=71
so 71/80 is your answer.